Swanson & Hiller, Inc., purchased a new machine on September 1 of the current year at a cost of $130,000. The machine’s estimated useful life at the time of the purchase was five years, and its residual value was $10,000. The company reports on a calendar year basis. Required:

a-1. Prepare a complete depreciation schedule, beginning with the current year, using the straight-line method. (Assume that the half-year convention is used).

a-2. Prepare a complete depreciation schedule, beginning with the current year, using the 200 percent declining-balance method. (Assume that the half-year convention is used).

a-3. Prepare a complete depreciation schedule, beginning with the current year, using the 150 percent declining-balance, switching to straight-line when that maximizes the expense.

(Assume that the half-year convention is used). b. Which of the three methods computed in part a is most common for financial reporting purposes?

c. Assume that Swanson & Hiller sells the machine on December 31 of the fourth year for $29,500 cash. Compute the resulting gain or loss from this sale under each of the depreciation methods used in part a.
Answer #1

a-1. Straight line method:

Depreciation expense = Purchase cost of asset - residual value / estimated useful life

= $130,000 - 10,000 / 5

= $24,000

Year Depreciation Accumulated depreciation
1( for 6 months) $12,000 $12,000
2 24,000 36,000
3 24,000 60,000
4 24,000 84,000
5 24,000 108,000
6 (for 6 months) 12,000 120,000
Total $120,000

a-2. 200% declining balance method:

Depreciation rate = 100 / 5 years * 200% = 40%

Year Book value at the start of the year Depreciation rate Depreciation Accumulated depreciation Book value at the end of the year
1( for 6 months) $130,000 20% $26,000 $26,000 $104,000
2 104,000 40% 41,600 67,600 62,400
3. 62,400 40% 24,960 92,560 37,440
4 37,440 40% 14,976 107,536 22,464
5 22,464 40% 8,986 116,522 13,478
6(for 6 months) 13,478 20% 2,696 119,218 10,782

a-3 150% declining balance method:

Depreciation rate = 100 / 5 years * 150% = 30%

Year Book value at the start of the year Depreciation rate Depreciation Accumulated depreciation Book value at the end of the year
1( for 6 months) $130,000 15% $19,500 $19,500 $110,500
2 110,500 30% 33,150 52,650 77,350
3. 77,350 30% 23,205 75,855 54,145
4 54,145 30% 16,244 92,099 37,901
5 37,901 Straight line method * 13,951 106,050 23,950
6 23,950 Straight line method 13,950 120,000 10,000

* Switch to straight line method from year 5 , Depreciation expense  calculated as (37,901 - 10,000 / 2) for both the remaining years

b. Straight line method of depreciation is the most common for financial reporting purposes out of the three methods.

c. Gain or( loss) from sale of machine = Sale value - Book value after depreciation ( cost of machine - accumulated depreciation)

Straight line method = $29,500 - 46,000 (130,000 - 84,000)

= $(16,500)

200% declining balance method = $29,500 - 22,464 (130,000 - 107,536)

= $7,036

150% declining balance method = $29,500 - 37,901 (130,000 - 92,099)

= $(8,401)